OpenStudy (anonymous):

Find the zeros of x^3-4x^2-5x=0 please help!!!

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (cwrw238):

first take out the common factor x: x(x^2 - 4x - 5) = 0

OpenStudy (anonymous):

Ok thaks what do I do next?

OpenStudy (cwrw238):

now can you factor x^2 - 4x - 5 ?

OpenStudy (anonymous):

No. I don't think so.

OpenStudy (cwrw238):

ok it will factor to 2 binomials of the form (x + a)(x + b) in this case a times b will = -5 and a+ b will = -4

OpenStudy (cwrw238):

so can you figure out what a and b are?

OpenStudy (cwrw238):

hint: a = -5

OpenStudy (anonymous):

(x-5) (x-4)?

OpenStudy (cwrw238):

no because -4 * -4 = -20 not -5

OpenStudy (cwrw238):

* -5 * -4 = 20 i meant!

OpenStudy (anonymous):

why would you do -5 * -4 though?

OpenStudy (cwrw238):

the last 2 numbers are multiplied to get the last number (x - 5)(x - 4) = x*x + x*-4 - 5* x -5*-4 = x^2 - 4x - 5x + 20 = x^2 - 9x + 20 which is nothing like what we require

OpenStudy (cwrw238):

first x is distributed over the (x - 4) then -5 is distributed over (x - 4)

OpenStudy (anonymous):

Oh oh so you FOIL'd it

OpenStudy (cwrw238):

yes - amounts to same thing the 2 numbers you require are - 5 and +1 because -5*+1 = -5 and -5+1 = -4

OpenStudy (anonymous):

Okay okay thank you. Is that all you do to find the zeros?

OpenStudy (cwrw238):

well you need to finish the factoring first: we have x(x - 5)(x + 1) = 0 so can u find the zeros now?

OpenStudy (anonymous):

Would x=5 and x=-1?

OpenStudy (cwrw238):

thats 2 zeros what about the other?

OpenStudy (anonymous):

I'm not sure what to do from there.